# VerAs: Verify then Assess STEM Lab Reports

### v1

@@ Abstract
With an increasing focus in STEM education on critical thinking skills, science writing plays an ever more important role in curricula that stress inquiry skills. A recently published dataset of two sets of college level lab reports from an inquiry-based physics curriculum relies on analytic assessment rubrics that utilize multiple dimensions, specifying subject matter knowledge and general components of good explanations. Each analytic dimension is assessed on a 6-point scale, to provide detailed feedback to students that can help them improve their science writing skills. Manual assessment can be slow, and difficult to calibrate for consistency across all students in large classes. While much work exists on automated assessment of open-ended questions in STEM subjects, there has been far less work on long-form writing such as lab reports. We present an end-to-end neural architecture that has separate verifier and assessment modules, inspired by approaches to Open Domain Question Answering (OpenQA). VerAs first verifies whether a report contains any content relevant to a given rubric dimension, and if so, assesses the relevant sentences. On the lab reports, VerAs outperforms multiple baselines based on OpenQA systems or Automated Essay Scoring (AES). VerAs also performs well on an analytic rubric for middle school physics essays.
@@
Panadero et al. [29] define a rubric as setting expectations for student work through specification of evaluative criteria, and how to meet the criteria. Their meta-review includes studies where rubrics do not necessarily have a scoring strategy, as when the main goal is formative assessment. Formative assessment occurs during a course or unit while students are learning the material, to help them improve before the end of the course. A rubric is analytic if it specifies multiple criteria, or rubric dimensions. We designed an automated approach for formative assessment and apply it to college level physics lab reports using a scoring strategy, and to middle school essays without one.
Fig. 1 illustrates one of the main challenges with the lab report rubrics: although all dimensions rely on a 6-point scale (0 to 5), they can use different scoring strategies. The top of the figure shows part of the first rubric dimension for lab reports on the behavior of a pendulum. The scoring strategy for this dimension is to determine which of five criteria are met, each with a different point value. The bottom of the figure shows part of the seventh dimension of a rubric for a report on Newton’s laws of force and motion. The scoring strategy enumerates the elements of a complete and correct discussion of random error, each with a point value, where the scoring strategy is to sum all the points. A second challenge is that in discursive science writing, it can be difficult to localize what part of a report is relevant for a given rubric dimension. As we discuss later in the paper, while human assessors can perform reliably on assigning a score for each dimension, they do not agree well on exactly which sentences address a given dimension.
| Pendulum D1: Is able the state the research question for reader clarity | |
@@
Given a rubric with dimensions and student lab reports, our assessment task is to generate a score for each dimension-report pair in the range . Inspired by Open-Domain Question Answering, we propose VerAs, which has a verifier module to determine whether a report contains sentences relevant to a dimension, and a grader to score the relevant sentences selected by the verifier. We test its effectiveness on a published dataset of lab reports [30] against multiple baselines. Through ablations, we demonstrate the need for both modules, and the benefit of using an ordinal loss training objective for the grader. We provide detailed error analysis of performance differences across rubric dimensions. To demonstrate the generality of the architecture, we also report results on middle school physics essays where the grader module is not necessary. We present related work, the datasets, VerAs architecture, experiments and results.
@@ 2 Related Work
Our approach is inspired by Open Domain Question Answering (OpenQA), where the goal is to find an answer to a question by querying multiple documents, some of which may contain no relevant information. Most OpenQA systems have two modules, a retriever to find relevant sentences, and a reader to extract the answer [16, 19, 18, 20, 24, 23, 37, 40]. Izacard & Grave [19] combine Dense Passage Retrieval (DPR) [20] with a sequence-to-sequence transformer reader module. In later work, they propose FiD-KD to perform knowledge distillation as a way to compensate for training data that lacks labeled pairs of queries and documents with answers [18]. Similarly, Read+Verify [17] has a distinct module to assess whether a question-passage pair can provide an answer. VerAs processes sentences rather than passages, but also relies on a verifier module to first determine whether a lab report contains sentences relevant to a given rubric dimension. Similar to [18, 17], we lack annotations on which sentences in a report, if any, are relevant to each rubric dimension.
@@ 3 Datasets
The college physics dataset consists of two sets of lab reports from a curriculum designed to promote scientific reasoning skills [30]. The first, about factors affecting the period of a pendulum, has a 7-dimension rubric. The second report, on Newton’s Second Law, has an 8-dimension rubric. Each rubric dimension specifies precise criteria for each point increment on a six-point scoring scale, as illustrated in Fig. 1; the appendix provides the complete rubrics. Each report has a ground truth score for each dimension from one of four trained raters. On random subsets of multiply labeled reports, raters had an average Pearson correlation of 0.72 on the 7 dimensions of the first report, and 0.69 on the 8 dimensions of the second report. The top half of Table 1 shows the size of the dataset splits (training, validation, test) and mean sentence lengths. As shown in Fig. 2, scores per dimension are highly skewed (cf. also Appendix 4).
The middle school data consists of responses to two essay prompts from a unit on the physics of roller coasters [33], as shown in the bottom half of Table 1. The first essay rubric identifies six main ideas about energy and the law of conservation of energy. The second essay rubric adds two additional ideas about the relations of mass to speed, and height to speed. Only 159 of the essays have reliable manual labels indicating the presence or absence of each main idea (Cohen’s kappa = 0.77) (essay 1 test is entirely manual labels). The remaining labels are from an automated tool called PyrEval [14, 41] whose accuracies on the two essays are 0.76 and 0.80, respectively, as reported below. For essay 2, there are reliable manual labels on 56 essays, corresponding to the essay 2 test set.
@@ 4.2 Grader
where is a linear layer, is the encoder, and , , and are the learned parameters of the encoders for the rubric dimension, report and prediction layers, respectively. We experiment with BERT [10], ELECTRA [8] and LongT5 [15] for the encoder function.
@@ 5 Experiments
All experiments use the Adam optimizer [21], and the same learning rates (0.001, 0.0001, 0.00001, 0.005, 0.0005, 0.00005) and batch sizes (4, 8, 16). We select the optimal hyperparameters given the validation loss, except for RBERT. Its loss automatically decreases each epoch because of its dynamic weight strategy, so we rely on the Spearman correlation instead. We tune for OLL with and . Lastly, we try and as the top parameter for both VerAs and FiD-KD. For FiD-KD, 25 is best, which is close to the average report length. For VerAs, 3 is best, which is close to the average of our two human raters (see section 4).
@@ 5.1 Baselines
Two of the baselines are distinct models, and one is a variant of VerAs. The RBERT [47] AES system predicts a total score given a report, without utilizing the rubric dimensions. It uses a BERT encoder followed by a linear layer to predict the score, scaling the scores to . The loss is a dynamically weighted sum of a regression (MSE) and ranking loss (CE). We tune the learning rate, batch size, and truncation size of the report.
@@
The third baseline reimplements VerAs as a multi-task model where each rubric dimension is treated as a separate problem, with a separate classification layer in the grader for each rubric dimension. The verifier module remains the same. VerAs thus tests whether different classifiers are needed to handle the semantic diversity across rubric dimensions.
@@ 5.2 Ablations
The first ablation replaces the verifier with random selection of three sentences from each lab report (Random Verifier). The second and third ablations omit the verifier module altogether, with the grader receiving only the rubric dimension and report, using either a truncated report to meet the input length constraint (W/o Verifier Trunc.), or an average of embeddings of a moving window over the full report (W/o Verifier Mov. Avg.). In the fourth ablation, the input to the grader omits the report (W/o Report). The final ablation, VerAs, uses cross entropy loss instead of OLL.
@@ 6.2 Results by Total Score and by Dimension
Table 2 shows that VerAs outperforms all of the baselines: by at least 17.9% on MSE, 8.0% on , and 0.8% on weighted accuracy. On total score, VerAs performs less well than VerAs, possibly because each classifier has only 1,666 examples instead of 12,460. Surprisingly, RBERT outperforms FiD-KD in two metrics although it uses a simpler architecture. VerAs also outperforms the ablations on MSE and , especially when CE instead of OLL is used. The weighted accuracy results are uniformly high due to the extreme data skew, but show no sensitivity across models. Table 3, which gives the average per dimension correlations and agreement, shows VerAs to have the highest performance, with VerAs outperforming FiD-KD.
@@ 6.3 Error Analysis of the Verifier’s Binary Decision
With respect to overall performance, Table 4 shows that the verifier does a better job on lab 1, which is also easier for the students: Fig. 2b) shows the mean score on lab 2 to be 2.16 in the training data, compared to 2.95 on lab 1, but with relatively few zero scores on any dimension (see Appendix). We speculate that the verifier does better on lab 1 because the data is more balanced. On each dimension, verifier accuracy is often close to the majority class baseline. However, for dimension 6 on lab 1, and dimensions 2-5 on lab 2, it is lower than the majority class result; for dimensions 1, 6 and 8 on lab 2, the verifier accuracy is greater than the majority class baseline. There appears to be a relationship between the difficulty of the rubric dimension and the performance of the verifier for the second lab. We calculate the pearson and spearman correlations between the accuracy of the verifier and the average training ground truth scores for each rubric dimension and we get 0.96 and 0.78 respectively. However, there is no such correlation for the first lab.
| Pendulum | F & M | ||||||||||
| Dim. | Maj. Base. | Acc. | Prec. | Rec. | F1 | Dim. | Maj. Base. | Acc. | Prec. | Rec. | F1 |
| 1 | 0.90 | 0.90 | 0.96 | 0.93 | 0.95 | 1 | 0.92 | 0.95 | 0.98 | 0.97 | 0.97 |
| 2 | 0.95 | 0.96 | 0.97 | 0.99 | 0.98 | 2 | 0.88 | 0.82 | 1.00 | 0.80 | 0.89 |
| 3 | 0.98 | 0.99 | 0.99 | 1.00 | 1.00 | 3 | 0.90 | 0.84 | 0.98 | 0.84 | 0.90 |
| 4 | 0.96 | 0.97 | 0.98 | 0.99 | 0.98 | 4 | 0.85 | 0.75 | 0.97 | 0.72 | 0.83 |
| 5 | 0.93 | 0.93 | 1.00 | 0.93 | 0.96 | 5 | 0.84 | 0.82 | 0.96 | 0.82 | 0.88 |
| 6 | 0.98 | 0.94 | 1.00 | 0.94 | 0.97 | 6 | 0.56 | 0.61 | 1.00 | 0.13 | 0.23 |
| 7 | 0.76 | 0.75 | 0.96 | 0.71 | 0.81 | 7 | 0.92 | 0.92 | 0.97 | 0.94 | 0.95 |
| 8 | 0.76 | 0.80 | 0.98 | 0.75 | 0.85 | ||||||
| Overall | 0.92 | 0.92 | 0.91 | 0.92 | 0.91 | Overall | 0.81 | 0.81 | 0.83 | 0.81 | 0.80 |
@@ 7 Conclusion
Our results show that formative assessment of longer forms of student writing, even those as complex as college-level lab reports with very detailed rubrics, can be handled by a neural network. VerAs performs very well on two sets of college level lab reports at applying a fine-grained analytic rubric, outperforming strong baselines. Ablations show that omitting the verifier module lowers MSE and on the total report score. This indicates the verifier plays an important role despite the lack of labeled data for the verifier sentence selection task. Evaluation of how well each dimension is scored, however, shows that VerAs outperforms VerAs. On a less complex essay dataset, VerAs, FiD-KD and a content assessment toolkit that requires no training perform equally well. Future work might focus on incorporating the score definitions in the rubrics and a better strategy to deal with the lack of labeled data for the sentence selection task.
Appendix 0.A Rubrics and Score Distributions

### v2

@@ Abstract
With an increasing focus in STEM education on critical thinking skills, science writing plays an ever more important role. A recently published dataset of two sets of college level lab reports from an inquiry-based physics curriculum relies on analytic assessment rubrics that utilize multiple dimensions, specifying subject matter knowledge and general components of good explanations. Each analytic dimension is assessed on a 6-point scale, to provide detailed feedback to students that can help them improve their science writing skills. Manual assessment can be slow, and difficult to calibrate for consistency across all students in large enrollment courses with many sections. While much work exists on automated assessment of open-ended questions in STEM subjects, there has been far less work on long-form writing such as lab reports. We present an end-to-end neural architecture that has separate verifier and assessment modules, inspired by approaches to Open Domain Question Answering (OpenQA). VerAs first verifies whether a report contains any content relevant to a given rubric dimension, and if so, assesses the relevant sentences. On the lab reports, VerAs outperforms multiple baselines based on OpenQA systems or Automated Essay Scoring (AES). VerAs also performs well on an analytic rubric for middle school physics essays.
@@
Panadero et al. [30] define a rubric as setting expectations for student work through specification of evaluative criteria, and how to meet them. Their meta-review includes studies where rubrics lack a scoring strategy, as when the main goal is formative assessment, which occurs during a course while students are learning the material, to help them improve by the end of the course. A rubric is analytic if it specifies multiple criteria, or rubric dimensions.We designed an automated approach for rubric assessment of STEM writing, and evaluated it on college level physics lab reports that have a scoring strategy in the rubric, and on middle school essays where there is no scoring strategy.
Fig. 1 illustrates a key challenge with the lab report rubrics: they can use different scoring strategies. The top of the figure shows part of the first rubric dimension for lab reports on the behavior of a pendulum. This is a criterion-based rubric where each point increment requires more explanation and correctness. The bottom of the figure shows part of the seventh dimension of a rubric for a report on Newton’s second law. Here an inclusion-based criterion is used, and the scoring strategy is to sum all the points.
A second challenge is that in discursive science writing, it can be difficult to localize what part of a report is relevant for a given rubric dimension. As we discuss later in the paper, while human assessors can perform reliably on assigning a score for each dimension, they do not agree well on exactly which sentences address a given dimension.
| Pendulum D1: Is able to state the research question for reader clarity | |
@@
Given a rubric with dimensions and student lab reports, our assessment task is to generate a score for each dimension-report pair in the range . Inspired by Open-Domain Question Answering, we propose VerAs111The code for VerAs is available at https://github.com/psunlpgroup/VerAs, which has a verifier module to determine whether a report contains sentences relevant to a dimension, and a grader to score the relevant sentences selected by the verifier. We test its effectiveness on a published dataset of lab reports [31] against multiple baselines. Through ablations, we demonstrate the need for both modules, and the benefit of using an ordinal loss training objective for the grader. We provide detailed error analysis of performance differences across rubric dimensions. To demonstrate the generality of the architecture, we also report results on middle school physics essays where the grader module is not necessary. We present related work, the datasets, VerAs architecture, experiments and results.
@@ 2 Related Work
Our approach is inspired by Open Domain Question Answering (OpenQA), where the goal is to query multiple documents, some of which may contain no relevant information. Most OpenQA systems have two modules, a retriever to find relevant sentences, and a reader to extract the answer [17, 19, 25, 24, 38, 41]. Izacard & Grave [20] combine Dense Passage Retrieval (DPR) [21] with a sequence-to-sequence transformer reader module. In later work, they propose FiD-KD to perform knowledge distillation as a way to compensate for training data that lacks labeled pairs of queries and documents with answers [19]. Similarly, Read+Verify [18] has a distinct module to assess whether a question-passage pair can provide an answer. VerAs processes sentences rather than passages, but also relies on a verifier module to first determine whether a lab report contains sentences relevant to a given rubric dimension. Similar to [19, 18], we lack annotations on which sentences in a report, if any, are relevant to each rubric dimension.
@@ 3 Datasets
The college physics dataset consists of two sets of lab reports [31] from a curriculum designed to promote scientific reasoning skills. The first, about factors affecting the period of a pendulum, has a 7-dimension rubric. The second report, on Newton’s Second Law, has an 8-dimension rubric. Each rubric dimension specifies precise criteria for each point increment on a six-point scoring scale, as illustrated in Fig. 1; the supplemental provides the complete rubrics. Each report has a ground truth score for each dimension from one of four trained raters. On random subsets of multiply labeled reports, raters had an average Pearson correlation of 0.72 on the 7 dimensions of the first report, and 0.69 on the 8 dimensions of the second report. The top half of Table 1 shows the size of the dataset splits (training, validation, test) and mean sentence lengths. As shown in Fig. 2, scores per dimension are highly skewed.
The middle school data consists of responses to two essay prompts from a unit on the physics of roller coasters [34], as shown in the bottom half of Table 1. The first essay rubric identifies six main ideas about energy and the law of conservation of energy. The second essay rubric adds two additional ideas about the relations of mass to speed, and height to speed. Only 159 of the essays have reliable manual labels indicating the presence of main ideas (Cohen’s kappa = 0.77) (essay 1 test is entirely manual labels). The remaining labels are from an automated tool called PyrEval [15, 42] whose accuracies on the two essays are 0.76 and 0.80, respectively, as reported below. For essay 2, there are reliable manual labels on 56 essays, corresponding to the essay 2 test set.
@@ 4.2 Grader
where is a linear layer, is the encoder, and , , and are the learned parameters of the encoders for the rubric dimension, report and prediction layers, respectively. We experiment with BERT [11] ELECTRA [9] and LongT5 [16] for the encoder function.
@@ 5 Experiments
All experiments use the Adam optimizer [22], and the same learning rates (0.001, 0.0001, 0.00001, 0.005, 0.0005, 0.00005) and batch sizes (4, 8, 16). We select the optimal hyperparameters given the validation loss, except for R2BERT. Its loss automatically decreases each epoch because of its dynamic weight strategy, so we rely on the Spearman correlation instead. We tune for OLL with and . Lastly, we try and as the top parameter for both VerAs and FiD-KD. For FiD-KD, 25 is best, which is close to the average report length. For VerAs, 3 is best, which is close to the average of our two human raters (see section 4).
@@ 5.1 Baselines
Two of the baselines are distinct models, and one is a variant of VerAs. The R2BERT [48] AES system predicts a total score given a report, without utilizing the rubric dimensions. It uses a BERT encoder followed by a linear layer to predict the score, scaling the scores to . The loss is a dynamically weighted sum of a regression (MSE) and ranking loss (CE). We tune the learning rate, batch size, and truncation size of the report.
@@
The third baseline reimplements VerAs as a multi-task model: each rubric dimension becomes a separate problem, with a separate classification layer in the grader for each rubric dimension. The verifier module remains the same. VerAsSEP thus tests whether different classifiers are needed to handle the semantic diversity across rubric dimensions.
@@ 5.2 Ablations
The first ablation replaces the verifier with random selection of three sentences from each lab report (Random Verifier). The second and third ablations omit the verifier module altogether, with the grader receiving only the rubric dimension and report, using either a truncated report to meet the input length constraint (W/o Verifier Trunc.), or an average of embeddings of a moving window over the full report (W/o Verifier Mov. Avg.). In the fourth ablation, the input to the grader omits the report (W/o Report). The final ablation, VerAsCE, uses cross entropy loss instead of OLL.
@@ 6.2 Results by Total Score and by Dimension
Table 2 shows that VerAs outperforms all of the baselines: by at least 17.9% on MSE, 8.0% on , and 0.8% on weighted accuracy. On total score, VerAsSEP performs less well than VerAs, possibly because each classifier has only 1,666 examples instead of 12,460. Surprisingly, R2BERT outperforms FiD-KD in two metrics although it uses a simpler architecture. VerAs also outperforms the ablations on MSE and , especially when CE instead of OLL is used. The weighted accuracy results are uniformly high due to the extreme data skew, but show no sensitivity across models. Table 3, which gives the average per dimension correlations and agreement, shows VerAsSEP to have the highest performance, with VerAs outperforming FiD-KD.
@@ 6.3 Error Analysis of the Verifier’s Binary Decision
With respect to overall performance, Table 4 shows that the verifier does a better job on lab 1, which is also easier for the students: Fig. 2b) shows the mean score on lab 2 to be 2.16 in the training data, compared to 2.95 on lab 1, but with relatively few zero scores on any dimension (see supplemental). We speculate that the verifier does better on lab 1 because the data is more balanced. On each dimension, verifier accuracy is often close to the majority class baseline. However, for dimension 6 on lab 1, and dimensions 2-5 on lab 2, it is lower than the majority class result; for dimensions 1, 6 and 8 on lab 2, the verifier accuracy is greater than the majority class baseline. In general, it provides good sentences, which is the more important responsibility of the verifier and through ablation studies, we show its effectiveness. There appears to be a relationship between the difficulty of the rubric dimension and the performance of the verifier for the second lab. We calculate the pearson and spearman correlations between the accuracy of the verifier and the average training ground truth scores for each rubric dimension and we get 0.96 and 0.78 respectively. However, there is no such correlation for the first lab.
| Pendulum | F & M | |||||||||||
| Dim. | Maj. Base. | Acc. | Prec. | Rec. | F1 | Dim. | Maj. Base. | Acc. | Prec. | Rec. | F1 | |
| 1 | 0.90 | 0.90 | 0.96 | 0.93 | 0.95 | 1 | 0.92 | 0.95 | 0.98 | 0.97 | 0.97 | |
| 2 | 0.95 | 0.96 | 0.97 | 0.99 | 0.98 | 2 | 0.88 | 0.82 | 1.00 | 0.80 | 0.89 | |
| 3 | 0.98 | 0.99 | 0.99 | 1.00 | 1.00 | 3 | 0.90 | 0.84 | 0.98 | 0.84 | 0.90 | |
| 4 | 0.96 | 0.97 | 0.98 | 0.99 | 0.98 | 4 | 0.85 | 0.75 | 0.97 | 0.72 | 0.83 | |
| 5 | 0.93 | 0.93 | 1.00 | 0.93 | 0.96 | 5 | 0.84 | 0.82 | 0.96 | 0.82 | 0.88 | |
| 6 | 0.98 | 0.94 | 1.00 | 0.94 | 0.97 | 6 | 0.56 | 0.61 | 1.00 | 0.13 | 0.23 | |
| 7 | 0.76 | 0.75 | 0.96 | 0.71 | 0.81 | 7 | 0.92 | 0.92 | 0.97 | 0.94 | 0.95 | |
| 8 | 0.76 | 0.80 | 0.98 | 0.75 | 0.85 | |||||||
| Overall | 0.92 | 0.92 | 0.91 | 0.92 | 0.91 | Overall | 0.81 | 0.81 | 0.83 | 0.81 | 0.80 | |
@@ 7 Conclusion
Our results show that formative assessment of longer forms of student writing, even those as complex as college-level lab reports with very detailed rubrics, can be handled by a neural network. VerAs performs very well on two sets of college level lab reports at applying a fine-grained analytic rubric, outperforming strong baselines. Ablations show that omitting the verifier module lowers MSE and on the total report score. This indicates the verifier plays an important role despite the lack of labeled data for the verifier sentence selection task. Evaluation of how well each dimension is scored, however, shows that VerAsSEP outperforms VerAs. On a less complex essay dataset, VerAs, FiD-KD and a content assessment toolkit that requires no training perform equally well. Future work might focus on incorporating the score definitions in the rubrics and a better strategy to deal with the lack of labeled data for the sentence selection task. Additionally, large language models such as GPT-4 [1] can be prompted to have potentially noisy labels for the relevant sentences.
Limitations: VerAs needs to be retrained for new datasets, which reduces its generality. Future work might focus on this by using large language models. We test VerAs performance on two college physics lab reports and one middle school physics essay. Future work might test on other STEM domains such as biology.
8 Acknowledgements
We thank Sarkar Das, Vipul Gupta, Zhaohui Li, and Ruihao Pan for helpful discussions. The second author’s work was supported by NSF DRK award 2010351.